# LoRA: Low-Rank Adaptation of Large Language Models

### v1

@@ Appendix F Measuring Similarity Between Subspaces
Appendix G Additional Experiments on Low-Rank Matrices
@@
G.1 Correlation between LoRA Modules
See Fig. 5 and Fig. 6 for how the results presented in Fig. 3 and Fig. 4 generalize to other layers.
@@ G.1 Correlation between LoRA Modules
G.2 Effect of $r$ on GPT-2
We repeat our experiment on the effect of $r$ (Sec. 6.2) in GPT-2. Using the E2E NLG Challenge dataset as an example, we report the validation loss and test metrics achieved by different choices of $r$ after training for 26,000 steps. We present our result in Table 14. The optimal rank for GPT-2 Medium is between 4 and 16 depending on the metric used, which is similar to that for GPT-3 175B. Note that the relationship between model size and the optimal rank for adaptation is still an open question.
@@ G.2 Effect of $r$ on GPT-2
G.3 Correlation between $W$ and $\Delta W$
See Fig. 7 for the normalized subspace similarity between $W$ and $\Delta W$ with varying $r$.
@@ G.4 Amplification Factor
Intuitively, when $\Delta W$ mostly contains task-specific directions, this quantity measures how much of them are amplified by $\Delta W$. As shown in Sec. 6.3, for $r=4$, this amplification factor is as large as 20. In other words, there are (generally speaking) four feature directions in each layer (out of the entire feature space from the pre-trained model $W$), that need to be amplified by a very large factor 20, in order to achieve our reported accuracy for the downstream specific task. And, one should expect a very different set of feature directions to be amplified for each different downstream task.

### v2

@@ Appendix G Measuring Similarity Between Subspaces
Appendix H Additional Experiments on Low-Rank Matrices
@@
H.1 Correlation between LoRA Modules
See Figure 6 and Figure 7 for how the results presented in Figure 3 and Figure 4 generalize to other layers.
@@ H.1 Correlation between LoRA Modules
H.2 Effect of $r$ on GPT-2
We repeat our experiment on the effect of $r$ (Section 7.2) in GPT-2. Using the E2E NLG Challenge dataset as an example, we report the validation loss and test metrics achieved by different choices of $r$ after training for 26,000 steps. We present our result in Table 18. The optimal rank for GPT-2 Medium is between 4 and 16 depending on the metric used, which is similar to that for GPT-3 175B. Note that the relationship between model size and the optimal rank for adaptation is still an open question.
@@ H.2 Effect of $r$ on GPT-2
H.3 Correlation between $W$ and $\Delta W$
See Figure 8 for the normalized subspace similarity between $W$ and $\Delta W$ with varying $r$.
@@ H.4 Amplification Factor
Intuitively, when $\Delta W$ mostly contains task-specific directions, this quantity measures how much of them are amplified by $\Delta W$. As shown in Section 7.3, for $r=4$, this amplification factor is as large as 20. In other words, there are (generally speaking) four feature directions in each layer (out of the entire feature space from the pre-trained model $W$), that need to be amplified by a very large factor 20, in order to achieve our reported accuracy for the downstream specific task. And, one should expect a very different set of feature directions to be amplified for each different downstream task.